1) Crazy stuntwoman, Ima Nutts, is performing a stunt for an upcoming summer movie thriller. In the stunt, she has to jump off the top of a very high apartment building and land in a crash mat below. At a point during the fall, when she is still 70.0 m above the crash mat, she is falling at a speed of 31 m/s. Her mass is 75 kg.

a) What is her gravitational potential energy at this point?

B) What is her kinetic energy at this point?

c) How much total energy does she have at this point?

d) How high is the apartment building?

e) What will her velocity be when she hits the crash mat?



1
Expert's answer
2021-07-15T10:56:08-0400

a)


"U=(70)(75)(9.8)=51450\ J"

b)


"K=0.5(75)(31)^2=36037.5\ J"


c)


"E=51450+36037.5=87487.5\ J"

d)


"87487.5=(75)(9.8)H\\H=119\ m"

e)


"87487.5=0.5(75)v^2\\v=48.3\frac{m}{s}"